42 A 2.3 ke object moving due north at 17 ms made contact with a barrier for 0.75 s. resulting io a linal velocity of 12 ms due east. What was the impulse on the ohjeet? MAGNITUDE OF IMPULS DIRECTION OF IMPUULSE 35 B of N 35° E of S 35 E of N 35 E of S A. 12 N-s B. 12 N S 48 N 48 N s D.
Answer #1

Impulse = change in momentum

Impulse = 2.3 (12 - 17)

Impulse = -11.5

Magnitude of Impulse = 11.5 Kg.m/s

Direction

theta = arctan (17 / 12)

theta = 54.78 degree North of east.

or

theta = 90 - 54.78 = 35.2 degree east of north

Therefore, correct option is A